A. 15.0 mL of nitrogen gas at 0.976 atm is compressed to a pressure of 1.18 atm. What is the new volume?

B. What is the pressure in 50.0 L gas cylinder that contains 5.50 grams of nitrogen, N2, at 25oC?
Answer #1

A)

Given:

Vi = 15.0 mL

Pi = 0.976 atm

Pf = 1.18 atm

use:

Pi*Vi = Pf*Vf

0.976 atm * 15.0 mL = 1.18 atm * Vf

Vf = 12.4 mL

Answer: 12.4 mL

5)

Molar mass of N2 = 28.02 g/mol

mass(N2)= 5.50 g

use:

number of mol of N2,

n = mass of N2/molar mass of N2

=(5.5 g)/(28.02 g/mol)

= 0.1963 mol

Given:

V = 50.0 L

n = 0.1963 mol

T = 25.0 oC

= (25.0+273) K

= 298 K

use:

P * V = n*R*T

P * 50 L = 0.1963 mol* 0.08206 atm.L/mol.K * 298 K

P = 0.0960 atm

Answer: 0.0960 atm